# Who Can Withstand Chat-Audio Attacks? An Evaluation Benchmark for Large Language Models

### v1

@@ Abstract
Adversarial audio attacks pose a significant threat to the growing use of large language models (LLMs) in voice-based human-machine interactions. While existing research has primarily focused on model-specific adversarial methods, real-world applications demand a more generalizable and universal approach to audio adversarial attacks. In this paper, we introduce the Chat-Audio Attacks (CAA) benchmark including four distinct types of audio attacks, which aims to explore the the vulnerabilities of LLMs to these audio attacks in conversational scenarios. To evaluate the robustness of LLMs, we propose three evaluation strategies: Standard Evaluation, utilizing traditional metrics to quantify model performance under attacks; GPT-4o-Based Evaluation, which simulates real-world conversational complexities; and Human Evaluation, offering insights into user perception and trust. We evaluate six state-of-the-art LLMs with voice interaction capabilities, including Gemini-1.5-Pro, GPT-4o, and others, using three distinct evaluation methods on the CAA benchmark. Our comprehensive analysis reveals the impact of four types of audio attacks on the performance of these models, demonstrating that GPT-4o exhibits the highest level of resilience. Our data can be accessed via the following link: CAA.
@@
An Evaluation Benchmark for Large Language Models
Wanqi Yang1∗, Yanda Li1∗, Meng Fang2, Yunchao Wei3, Tianyi Zhou4, Ling Chen1 1University of Technology Sydney, 2University of Liverpool, 3Beijing Jiaotong University, 4Agency for Science, Technology and Research (A*STAR), Singapore {wanqi.yang-1, yanda.li}@student.uts.edu.au
11footnotetext: Equal contribution
@@ 1 Introduction
Large language models (LLMs) capable of processing text, images, and audio have become increasingly essential for applications that require advanced comprehension and response generation, including customer service Kolasani (2023); Hadi et al. (2024), automated content creation Todd et al. (2023); Sudhakaran et al. (2024), and interactive media He et al. (2023); Köpf et al. (2024). However, the versatile capabilities of these models also increase their vulnerability to adversarial attacks Shayegani et al. (2023); Zhao et al. (2024). This is particularly true in the domain of LLM-driven human-machine voice interaction, where the emergence of such services has accelerated research into audio-based adversarial attacks and defense mechanisms.
@@
Attacks on multimodal audio LLMs can cause the models to produce unintended outputs. However, this area has received limited attention, primarily due to the challenges associated with audio as an input modality. Unlike images, audio lacks direct gradient signals, making the crafting of adversarial examples more complex. Previous research on adversarial audio attacks has predominantly focused on targeted attacks Gong and Poellabauer (2017); Kassis and Hengartner (2021); Zhang et al. (2022), where carefully crafted perturbations are embedded within speech signals. While these samples can effective in misleading models, often appear as random noise and are easily detectable by human listeners. A notable advancement Carlini and Wagner (2018) introduced a gradient-based optimization approach that utilizes the Connectionist Temporal Classification (CTC) loss Graves et al. (2006)—a method designed for time series data in classification tasks. However, this method remains model-specific and lacks broader generalizability. Universal adversarial audio attacks Xie et al. (2021) are highly relevant to real-world attack scenarios, such as when a speaker makes a verbal error or when they are speaking in a noisy environment. Attackers can pre-design and generate these universal attacks in advance, then apply them to any input audio. Despite their relevance, there has been insufficient exploration of their impact on multimodal audio LLMs.
@@
As multimodal audio LLMs become more prevalent in human-machine voice interactions, the threat posed by these attacks grows significantly. To explore the vulnerabilities of LLMs to adversarial audio attacks, we propose a benchmark of universal adversarial audio attacks specifically based on conversational scenarios, named Chat-Audio Attacks (CAA). The CAA benchmark consists of 360 adversarial audio attack sets, with each set encompassing four distinct types of audio attacks: content attack, emotional attack, explicit noise attack, and implicit noise attack. This results in a total of 1,680 adversarial audio samples. We believe that CAA benchmark will not only enable researchers to pinpoint weaknesses in LLMs under adversarial audio conditions but also drive the advancement of robust defense mechanisms for LLM.
@@
In addition, we introduce three evaluation methods to comprehensively assess the resilience of LLMs against adversarial audio attacks: Standard Evaluation, GPT-4o-Based Evaluation, and Human Evaluation. The Standard Evaluation uses rigorous metrics to quantify the accuracy, similarity, and consistency of voice responses under adversarial conditions, providing a repeatable and controlled result. In contrast, the GPT-4o-Based Evaluation simulates real-world interactions, capturing complex, sensitive inaccuracies that standard metrics might overlook. Human Evaluation reflects actual user experience and perceptions, thus offering crucial insights into user trust.
Finally, we evaluate six state-of-the-art LLMs supporting voice-based conversations on the CAA benchmark, such as Gemini-1.5-Pro Reid et al. (2024) and GPT-4o OpenAI (2023), providing results across the three aforementioned evaluation methods. We analyzed the impact of four types of audio attacks on the LLMs and discussed the flaws these models exhibit in the face of such attacks.
@@
  • We propose three evaluation methods to systematically evaluate the performance of LLMs against adversarial audio attacks.
  • We perform a comprehensive evaluation of six state-of-the-art LLMs using the CAA benchmark. Based on the three experimental results, we provide an in-depth analysis and discussion of the results.
@@ 2 CAA Benchmark
In CAA benchmark, we target the response generation task by collecting audio suitable for human-machine chat. The overview of CAA benchmark is shown in Figure 1. Each set of audio attack data consists of a quadruplet ($a_{i}$, $t_{i}$, ${a_{i}}^{no\_attack}$ $\mathcal{A}_{i}$), where $a_{i}$ represents the original, unprocessed audio containing a single utterance; $t_{i}$ refers to the transcript of the original audio along with other associated textual labels; ${a_{i}}^{no\_attack}$ indicates the audio generated by a voice agent reading the transcript without any attack; and the set $\mathcal{A}_{i}$ includes 3 or 5 types of attack variations of the audio.
@@ 2.1 Audio Collection
For the unprocessed audio $a_{i}$ and corresponding transcripts $t_{i}$ in CAA benchmark, we manually collected data from three publicly available multimodal datasets (text, audio, and visual): MELD, TVQA, and Common Voice.
@@
  • Common Voice Ardila et al. (2019): is a multilingual dataset for speech recognition, it provides audios and transcripts. However, the audio samples are not explicitly designed in a dialogue format.
@@ 2.2 Audio Attack Generation
We observed that some samples from MELD, TVQA, and Common Voice are often impacted by factors such as speech rate, accent, and clarity, which can obscure the audio information, making them unsuitable as baselines for subsequent comparison and analysis. To address this, we generated no-attack audio to ensure that the LLMs receive clear speech inputs. This serves as a baseline, offering audio free from any interference or alterations.
@@
The goal of content attacks is to explore whether LLMs are sensitive to token changes or minor errors when the overall meaning of the audio remains preserved.
Emotional Attack alters the emotional tone of the audio without changing the content. CAA benchmark contains two types of emotional attacks: (1) opposing emotional tone, and (2) opposing emotional background music. In the first scenario, the AzureSpeechSDK agent was instructed to re-read the transcript with an emotion opposite to the original. For instance, if the original sample had an “angry” emotion label, the agent would re-read the transcripts with a “happy” tone, generating an opposite-emotion audio sample. In the second scenario, we overlaid background music with an opposing emotion onto the no-attack audio, and adjusted the music volume to ensure the speaker’s voice remained clear. It is important to emphasize that only samples from the MELD dataset in our collection are labeled with emotions. As a result, we utilized 120 samples from MELD to generate two emotional attack audio samples for each, resulting in two distinct emotional tones per sample.
The objective of emotional attacks is to investigate the sensitivity of LLMs to variations in speech emotion and whether a mismatch between speech content and emotional tone influences the responses.
@@
Explicit noise attacks is to evaluate the ability of LLMs to differentiate between the speaker’s voice and background noise, as well as to assess their robustness to such interference.
@@
The objective of implicit noise attacks is to assess whether LLMs, similar to humans, remain unaffected by inaudible noise.
@@ 3.1 Experimental Setup
Models We present a comprehensive performance evaluation of the most popular multimodal audio models, including SpeechGPT Zhang et al. (2023), SALMONN Tang et al. (2023), Qwen2-Audio Chu et al. (2024), LLama-Omni Fang et al. (2024), Gemini-1.5-pro Reid et al. (2024) and GPT-4o Achiam et al. (2023).
@@
Table 2: Standard evaluation results on CAA benchmark. Performance comparison of the multimodal audio LLMs under various adversarial conditions using WER, ROUGE-L, and COS metrics.
@@
Table 3: GPT-4o-based evaluation results on CAA benchmark. Performance comparison of the multimodal audio LLMs under various adversarial conditions using NC, ACoh, ACor and LR metrics.
@@ 3.2 Standard Evaluation
As shown in table 2, presenting the performance of the models across these metrics, comparing how well they handle adversarial interference in the audio inputs.
@@ 4 Discussion
Whether LLMs are sensitive to token changes or minor errors?
It is evident that different LLMs exhibit varying degrees of sensitivity to token changes or minor errors. GPT-4o consistently shows strong robustness across most metrics (WER, ROUGE-L, COS, ACoh, ACor, and LR), indicating lower sensitivity to token-level adversarial attack. In contrast, SpeechGPT and Qwen2-Audio exhibit greater vulnerability, with lower scores in these key areas, suggesting that minor token changes can significantly degrade their performance.
Is it good news that LLMs are unaffected by the mismatch between speech content and emotional tone?
We argue that it is not good news that LLMs remain unaffected by emotional mismatches. Although large language models demonstrate resilience by maintaining high levels of coherence, correlation, and semantic similarity, this also reflects their relative weakness in emotional awareness. Current LLMs still have considerable scope for improvement in recognizing emotional subtleties, as humans can easily detect emotional mismatches, such as sarcasm or passive-aggressive tones in conversations. While SpeechGPT is notably impacted by mismatches between speech content and emotional tone, this does not indicate a heightened sensitivity to emotional shifts, as its overall coherence score remains relatively low.
Which explicit noise attacks have the most significant impact on LLMs?
Natural noise has the most significant overall impact on LLMs across all metrics, especially on SpeechGPT and SALMONN, which shows the highest sensitivity to it. Industrial noise also causes notable attacks but is handled better by LLMs like Llama-Omni and Gemini-1.5-Pro. Human noise, while still impactful, is generally less detrimental compared to the other explicit noises. Overall, SpeechGPT and SALMONN show the most vulnerability across all types of explicit noise attacks, while Llama-Omni, Gemini-1.5-Pro and GPT-4o demonstrate stronger robustness.
Whether LLMs remain unaffected by inaudible noise?
@@
In summary, extensive data-driven pre-training appear to be key factors in helping models like GPT-4o stay robust against adversarial audio. This element enables the models to handle a variety of adversarial scenarios with minimal degradation in performance.
@@ 5.1 Audio/Speech Language Models
In the field of audio-based language models, initial systems Lakhotia et al. (2021); Radford et al. (2023); Borsos et al. (2022) utilized either acoustic or semantic tokens to enable generation from audio inputs into text or audio outputs. With the technological advancements brought by large language models (LLMs), the recent trend has shifted towards multimodal models Tang et al. (2023); Chen et al. (2023); Wu et al. (2023); Fathullah et al. (2024) are leveraging the combined strengths of both speech and text modalities, substantially enhancing the versatility and effectiveness of audio-based applications.
@@ 5.2 Audio Attacks
The initial generation of adversarial samples utilized optimization methods first developed for music genre classification Kereliuk et al. (2015). These techniques manipulated entire audio waveforms to avoid detection, altering not only specific acoustic features but the entire sound profile while preserving perceptual quality. In contrast, in the field of speech paralinguistics Gong and Poellabauer (2017); Kassis and Hengartner (2021); Zhang et al. (2022), the Fast Gradient Sign Method (FGSM) has been employed to craft adversarial samples aimed at disrupting systems. Large language models (LLMs) that process diverse data types such as text, images, and audio offer enhanced capabilities for generating human-like responses across various applications. However, their multi-modal nature also increases vulnerability to jailbreaks aud (2023) and adversarial attacks, with potential exploits spanning across all processed modalities, allowing attackers to bypass safety constraints embedded within these models.

### v2

# Who Can Withstand Chat-Audio Attacks? An Evaluation Benchmark for Large Audio-Language Models

@@ Abstract
Adversarial audio attacks pose a significant threat to the growing use of large audio-language models (LALMs) in voice-based human-machine interactions. While existing research focused on model-specific adversarial methods, real-world applications demand a more generalizable and universal approach to audio adversarial attacks. In this paper, we introduce the Chat-Audio Attacks (CAA) benchmark including four distinct types of audio attacks, which aims to explore the vulnerabilities of LALMs to these audio attacks in conversational scenarios. To evaluate the robustness of LALMs, we propose three evaluation strategies: Standard Evaluation, utilizing traditional metrics to quantify model performance under attacks; GPT-4o-Based Evaluation, which simulates real-world conversational complexities; and Human Evaluation, offering insights into user perception and trust. We evaluate six state-of-the-art LALMs with voice interaction capabilities, including Gemini-1.5-Pro, GPT-4o, and others, using three distinct evaluation methods on the CAA benchmark. Our comprehensive analysis reveals the impact of four types of audio attacks on the performance of these models, demonstrating that GPT-4o exhibits the highest level of resilience. Our data can be accessed via the following link: CAA.
@@
An Evaluation Benchmark for Large Audio-Language Models
Wanqi Yang1∗,  Yanda Li1thanks: Equal contributions,  Meng Fang2,  Yunchao Wei3,  Ling Chen1 1 University of Technology Sydney, 2 University of Liverpool, 3 Beijing Jiaotong University wanqi.yang-1@student.uts.edu.au, Yanda.Li@student.uts.edu.au Meng.Fang@liverpool.ac.uk, wychao1987@gmail.com, ling.chen@uts.edu.au
@@ 1 Introduction
Large language models (LLMs) capable of processing text, images, and audio have become increasingly essential for applications that require advanced comprehension and response generation, including customer service Kolasani (2023); Hadi et al. (2024), automated content creation Todd et al. (2023); Sudhakaran et al. (2024), and conversational systems He et al. (2023); Köpf et al. (2024); yang2025mtpchat. However, the versatile capabilities of these models also increase their vulnerability to adversarial attacks Shayegani et al. (2023); Zhao et al. (2024). This is particularly true in the domain of LLM-driven human-machine voice interaction, where the emergence of such services has accelerated research into audio-based adversarial attacks and defense mechanisms.
@@
Attacks on large audio-language models (LALMs) can cause the models to produce unintended outputs. However, this area has received limited attention, primarily due to the challenges associated with audio as an input modality. Unlike images, audio lacks direct gradient signals, making the crafting of adversarial examples more complex. Previous research on adversarial audio attacks has focused primarily on targeted attacks Gong and Poellabauer (2017); Kassis and Hengartner (2021); Zhang et al. (2022), where carefully crafted perturbations are embedded within speech signals. While these samples are effective in misleading models, they often appear as random noise and are easily detectable by human listeners. A notable advancement Carlini and Wagner (2018) introduced a gradient-based optimization approach that utilizes the Connectionist Temporal Classification loss Graves et al. (2006)—a method designed for time series data in classification tasks. However, this method remains model-specific and lacks broader generalizability. Universal adversarial audio attacks Xie et al. (2021) are highly relevant to real-world attack scenarios, such as when a speaker makes a verbal error or when they are speaking in a noisy environment. Attackers can pre-design and generate these universal attacks in advance, then apply them to any input audio. Despite their relevance, there has been insufficient exploration of their impact on LALMs.
@@
As LALMs become more prevalent in human-machine voice interactions, the threat posed by these attacks grows significantly. To explore the vulnerabilities of LALMs to adversarial audio attacks, we propose a benchmark of universal adversarial audio attacks specifically based on conversational scenarios, named Chat-Audio Attacks (CAA). The CAA benchmark consists of 360 adversarial audio attack sets, with each set encompassing four distinct types of audio attacks: content attack, emotional attack, explicit noise attack, and implicit noise attack. This results in a total of 1,680 adversarial audio samples. We believe that CAA benchmark will not only enable researchers to pinpoint weaknesses in LALMs under adversarial audio conditions but also drive the advancement of robust defense mechanisms for LALMs.
@@
In addition, we introduce three evaluation methods to comprehensively assess the resilience of LALMs against adversarial audio attacks: Standard Evaluation, GPT-4o-Based Evaluation, and Human Evaluation. The Standard Evaluation uses rigorous metrics to quantify the accuracy, similarity, and consistency of voice responses under adversarial conditions, providing a repeatable and controlled result. In contrast, the GPT-4o-Based Evaluation simulates real-world interactions, capturing complex, sensitive inaccuracies that standard metrics might overlook. Human Evaluation reflects actual user experience and perceptions, offering crucial insights into user trust.
Finally, we evaluate six state-of-the-art LALMs supporting voice-based conversations on the CAA benchmark, such as Gemini-1.5-Pro Reid et al. (2024) and GPT-4o OpenAI (2023), providing results across the three aforementioned evaluation methods. We analyze the impact of four types of audio attacks on the LALMs and discusse the flaws these models exhibit in the face of such attacks.
@@
  • We propose three evaluation methods to systematically evaluate the performance of LALMs against adversarial audio attacks.
  • We perform a comprehensive evaluation of six state-of-the-art LALMs using the CAA benchmark. Based on the three experimental results, we provide an in-depth analysis and discussion of the results.
@@ 2 CAA Benchmark
In CAA benchmark, we target the response generation task by collecting suitable audio for human-machine chat. The overview of CAA benchmark is shown in Figure 1. Each set of audio attack data consists of a quadruplet ($a_{i}$, $t_{i}$, ${a_{i}}^{no\_attack}$ $\mathcal{A}_{i}$), where $a_{i}$ represents the original, unprocessed audio containing a single utterance; $t_{i}$ refers to the transcript of the original audio along with other associated textual labels; ${a_{i}}^{no\_attack}$ indicates the audio generated by a voice agent reading the transcript without any attack; and the set $\mathcal{A}_{i}$ includes 3 or 5 types of attack variations of the audio.
@@ 2.1 Audio Collection
For unprocessed audio $a_{i}$ and corresponding transcripts $t_{i}$ in CAA benchmark, we manually collected data from three publicly available multimodal datasets (text, audio, and visual): MELD, TVQA, and Common Voice.
@@
  • Common Voice Ardila et al. (2019): is a multilingual dataset for speech recognition, provides audios and transcripts. However, the audio samples are not explicitly designed in a dialogue format.
@@ 2.2 Audio Attack Generation
We observed that some samples from MELD, TVQA, and Common Voice are often impacted by factors such as speech rate, accent, and clarity, which can obscure the audio information, making them unsuitable as baselines for subsequent comparison and analysis. To address this, we generated no-attack audio to ensure that the LALMs receive clear speech inputs. This serves as a baseline, offering audio free from interference or alterations.
@@
The goal of content attacks is to explore whether LALMs are sensitive to token changes or minor errors when the overall meaning of the audio remains preserved.
Emotional Attack alters the emotional tone of the audio without changing the content. CAA benchmark contains two types of emotional attacks: (1) opposing emotional tone, and (2) opposing emotional background music. In the first scenario, the AzureSpeechSDK agent was instructed to re-read the transcript with an emotion opposite to the original. For instance, if the original sample had an “angry” emotion label, the agent would re-read the transcripts with a “happy” tone, generating an opposite-emotion audio sample. In the second scenario, we overlaid background music with an opposing emotion onto the no-attack audio and adjusted the music volume to ensure the speaker’s voice remained clear. It is important to emphasize that only samples from the MELD dataset in our collection are labeled with emotions. As a result, we utilized 120 samples from MELD to generate two emotional attack audio samples for each, resulting in two distinct emotional tones per sample.
The objective of emotional attacks is to investigate the sensitivity of LALMs to variations in speech emotion and whether a mismatch between speech content and emotional tone influences responses.
@@
Explicit noise attacks are used to evaluate the ability of LALMs to differentiate between the speaker’s voice and background noise, as well as to assess their robustness to such interference.
@@
The objective of implicit noise attacks is to assess whether LALMs, similar to humans, remain unaffected by inaudible noise.
@@ 3.1 Experimental Setup
Models We present a comprehensive performance evaluation of the most popular large audio-language models, including SpeechGPT Zhang et al. (2023), SALMONN Tang et al. (2023), Qwen2-Audio Chu et al. (2024), LLama-Omni Fang et al. (2024), Gemini-1.5-pro Reid et al. (2024) and GPT-4o Achiam et al. (2023).
@@
Table 2: Standard evaluation results on CAA benchmark. Performance comparison of the LALMs under various adversarial conditions using WER, ROUGE-L, and COS metrics.
@@
Table 3: GPT-4o-based evaluation results on CAA benchmark. Performance comparison of the LALMs under various adversarial conditions using NC, ACoh, ACor and LR metrics.
@@ 3.2 Standard Evaluation
As shown in Table 2, presenting the performance of the models across these metrics and comparing how well they handle adversarial interference on the audio inputs.
@@ 4 Discussion
Are LALMs sensitive to token changes or minor errors?
It is evident that different LALMs exhibit varying degrees of sensitivity to token changes or minor errors. GPT-4o consistently shows strong robustness across most metrics (WER, ROUGE-L, COS, ACoh, ACor, and LR), indicating lower sensitivity to token-level adversarial attack. In contrast, SpeechGPT and Qwen2-Audio exhibit greater vulnerability, with lower scores in these key areas, suggesting that minor token changes can significantly degrade their performance.
Is it good news that LALMs are unaffected by the mismatch between speech content and emotional tone?
We argue that it is not good news that LALMs remain unaffected by emotional mismatches. Although large language models demonstrate resilience by maintaining high levels of coherence, correlation, and semantic similarity, this also reflects their relative weakness in emotional awareness. Current LALMs still have considerable scope for improvement in recognizing emotional subtleties, as humans can easily detect emotional mismatches, such as sarcasm or passive-aggressive tones in conversations. While SpeechGPT is notably impacted by mismatches between speech content and emotional tone, this does not indicate a heightened sensitivity to emotional shifts, as its overall coherence score remains relatively low.
Which explicit noise attacks have the most significant impact on LALMs?
Natural noise has the most significant overall impact on LALMs across all metrics, especially on SpeechGPT and SALMONN, which shows the highest sensitivity to it. Industrial noise also causes notable attacks but is handled better by LALMs like Llama-Omni and Gemini-1.5-Pro. Human noise, while still impactful, is generally less detrimental compared to the other explicit noises. Overall, SpeechGPT and SALMONN show the most vulnerability across all types of explicit noise attacks, while Llama-Omni, Gemini-1.5-Pro and GPT-4o demonstrate stronger robustness.
Are LALMs unaffected by inaudible noise?
@@
How architectural or training differences contribute to robustness?
1) Vulnerability of Models with Transcription Modules: Models such as SpeechGPT, which incorporate transcription modules, are particularly fragile. These models first transcribe input audio into text before performing inference. Consequently, the accuracy of the transcription process becomes critical. Our adversarial samples significantly degrade the performance of these transcription modules, leading to substantial drops in transcription accuracy when processing audio with noise.
2) Training Limitations of SALMONN: SALMONN lacks a task specifically designed for generating responses during training. As a result, both human and automated evaluations rated SALMONN’s responses poorly in terms of satisfaction. This shortcoming causes the model to behave inconsistently under attack: at times transcribing the audio, other times describing its content, and occasionally speculating about the speaker’s emotions. This inconsistency further undermines its robustness.
3) Robustness of Qwen2-Audio and Llama-Omni: Unlike SpeechGPT and SALMONN, Qwen2-Audio and Llama-Omni do not rely on transcribing input audio into text. Moreover, their training datasets include audio with noise, such as laughter, which contributes to their superior robustness against adversarial attacks.
4) Observations on GPT-4o and Gemini-1.5-Pro: While GPT-4o and Gemini-1.5-Pro have not released their model codes, an intriguing observation is their heightened vulnerability to infrasound and ultrasound attacks. In contrast, SALMONN is minimally affected by such attacks. This highlights an area for improvement in these models.
How these findings might influence model design or inform strategies for improving adversarial robustness?
Based on the analysis of the model structure and training methods presented above, we propose several directions for future research on speech models:
1) Developing enhanced transcription modules that demonstrate greater robustness to audio perturbations, or exploring alternative approaches that eliminate the reliance on transcription altogether.
2) Incorporating multi-task training objectives to improve the adaptability and generalizability of speech models.
3) Diversifying training datasets to better reflect real-world scenarios and adversarial conditions, ensuring broader applicability and resilience.
4) Designing targeted defense mechanisms to counter specific adversarial techniques, such as infrasound and ultrasound attacks. These directions aim to address existing challenges while paving the way for more robust and versatile speech models.
@@ 5.1 Audio/Speech Language Models
In the field of LALMs, initial systems Lakhotia et al. (2021); Radford et al. (2023); Borsos et al. (2022) utilized either acoustic or semantic tokens to enable generation from audio inputs into text or audio outputs. With the technological advancements brought by LLMs, the recent trend has shifted towards multimodal models Tang et al. (2023); Chen et al. (2023); Wu et al. (2023); Fathullah et al. (2024) are leveraging the combined strengths of both speech and text modalities, substantially enhancing the versatility and effectiveness of audio-based applications.
@@ 5.2 Audio Attacks
The initial generation of adversarial samples utilized optimization methods first developed for music genre classification Kereliuk et al. (2015). These techniques manipulated entire audio waveforms to avoid detection, altering not only specific acoustic features but the entire sound profile while preserving perceptual quality. In contrast, in the field of speech paralinguistics Gong and Poellabauer (2017); Kassis and Hengartner (2021); Zhang et al. (2022), the Fast Gradient Sign Method has been employed to craft adversarial samples aimed at disrupting systems. LLMs that process diverse data types such as text, images, and audio offer enhanced capabilities for generating human-like responses across various applications. However, their multi-modal nature also increases vulnerability to jailbreaks aud (2023) and adversarial attacks, with potential exploits spanning across all processed modalities, allowing attackers to bypass safety constraints embedded within these models.